OpenStudy (anonymous):

Substitutinginto Equations g(5)=27. Can any body please explain how the answer is 27. Thank you.

OpenStudy (anonymous):

eqn????

OpenStudy (anonymous):

very much

hartnn (hartnn):

where is g(x) ? is it x^2+2 ??

OpenStudy (anonymous):

Not sure

hartnn (hartnn):

there must be definations of g(x),f(x) and h(x) somewhere, search it.

OpenStudy (anonymous):

where is the fn G(x)???

OpenStudy (anonymous):

g(x) = function which consists of 'x' as a variable, if u put g(5) then replace x with 5 and hence you will get the desired answer :)

OpenStudy (anonymous):

so what is the equation really in the format: g(x)=__________ You have to tell first what the equation is. But whatever the equation is, you just substitute the value of x given to you which is 5 and then solve the equation from there. This will end up giving you the answer g(5)=27

OpenStudy (anonymous):

The question does not provide me with the equation.

OpenStudy (anonymous):

so you have to figure out the equation. Is that what you are saying???

OpenStudy (anonymous):

It seems like it.

OpenStudy (anonymous):

that is hard. there could be several possible equations for this solution. how can you figure out which one the question asks for?

OpenStudy (anonymous):

It maybe a typo, I will check with my teacher monday. Thanks for the help anyways.

OpenStudy (anonymous):

for example: one possible equation could be: g(x)=5x+2 so plugging the value of x that is 5: g(5)=5(5)+2 g(5)=25+2 g(5)=27 this is only one possible solution you can work out many of them

OpenStudy (anonymous):

got it, Thank you!

OpenStudy (anonymous):

@Bhagyashree if its g(x) = x^2 + 2

OpenStudy (anonymous):

that works too I just said there are many possible solutions or equations that you can work out from the given information